Darren is three years older than his sister.In three years, he will be twice as old as she will be then.How old is each now?














Answer by stanbon(75887)

Darren is three years older than his sister.In three years, he will be twice as old as she will be then.How old is each now?
--------------------------------
Let Darren be "d" now.
Let Sister be "s" now.
-----------------------------
Equations:
d = s + 3
(d+3) = 2(s+3)
----
Substitute for "d" and solve for "s":
(s+3+3) = 2(s+3)
s+6 = 2s+6
---
s = 0 (sister's age now)
---
d = s+3
d = 3 yrs old (Darren's age now)
===========================
Cheers,
Stan H.
==========